OpenStudy (anonymous):

Bart bought a digital camera with a list price of $219 from an online store offering a 6 percent discount. He needs to pay $7.50 for shipping. What was Bart's total cost?

OpenStudy (campbell_st):

Cost = camera - discount + shipping = 219 - 6/100 x 219 + 7.50 just do the calculations for the cost

OpenStudy (anonymous):

The manufacturing cost of a cell phone was $136. Sellers first marked up the cost by 25 percent. Sales were poor, so sellers decreased the markup to 17 percent. Calculate the old and new selling prices.

OpenStudy (campbell_st):

old selling price = 136 + 25/100 x 136 new selling price = old selling price - 17/100 x old selling price.

OpenStudy (anonymous):

Percy paid $24.10 for a basketball. The price of the basketball was $22.99. What was the sales tax rate?

OpenStudy (campbell_st):

sorry thats enough for me... good luck with your percentages work.

OpenStudy (anonymous):

No, Thank you :)